# The Color-Clinical Decoupling: Why Perceptual Calibration Fails Clinical Biomarkers in Smartphone Dermatology

### v1

@@ 5.3 Implications for Tele-Dermatology Standards
Third, the statistical significance of gender effects in our ANOVA ($p<0.001$) suggests that camera ISPs may exhibit differential behavior across demographics. If “universal” CCMs are derived from specific populations, they may embed demographic biases—a fairness concern for global health applications.[Kinyanjui2020, Groh2021]
@@ 6 Conclusion
This study establishes that standard colorimetric calibration in smartphone dermatology creates false confidence: perceptually accurate images can yield unreliable clinical biomarkers. Using 43,425 images from 965 Korean subjects across DSLR, tablet, and smartphone devices, we demonstrate a fundamental “color-clinical decoupling” phenomenon.

### v2

@@ 5.3 Implications for Tele-Dermatology Standards
Third, interestingly, gender also showed a statistically significant difference ($p<0.001$). Although its influence ($\eta^{2}=0.007$) is small compared to facial region, this suggests that camera ISPs may exhibit differential behavior across demographics. If “universal” CCMs are derived from specific populations, they may embed demographic biases—a fairness concern for global health applications.[Kinyanjui2020, Groh2021]
@@ 6 Conclusion
This study establishes that standard colorimetric calibration in smartphone dermatology creates false confidence: perceptually accurate images can yield unreliable clinical biomarkers. Using 43,424 images from 965 Korean subjects across DSLR, tablet, and smartphone devices, we demonstrate a fundamental “color-clinical decoupling” phenomenon.
@@
  3. 3.
    Calibration errors disproportionately affect darker skin: ITA sensitivity to $b^{*}$ errors increases from $\sim$1.7°/unit for light skin (Type II/III) to >>4.5°/unit for dark skin (Type VI). A 2-unit calibration residual produces a 3.4° ITA shift in light skin but a 9.2° shift in dark skin—sufficient to cause skin type misclassification. This “metrological bias” means current calibration protocols fail to protect the most vulnerable populations from measurement error.